OpenStudy (anonymous):

y varies directly with x, and y=2 when x=5. what is the value of x when y=15

OpenStudy (texaschic101):

5/2 = x/15 (cross multiply) 2x = (15)(5) Now just solve for x and you will have your answer Can you take it from here ?